Find the missing coefficient so that lines are perpendicular. x-3y= 2 and y=Bx+17
Answer by Alan3354(69443)

Find the missing coefficient so that lines are perpendicular. x-3y= 2 and y=Bx+17
---------------
Find the slope m of x-3y = 2
B = -1/m